Consider the following market. There are two contracts A and B available. Contract A is priced...

  1. Consider the following market. There are two contracts A and B available. Contract A is priced at $16 today and it delivers 2 gallons of oil and 4 bushels of corn tomorrow. Contract B is priced at $12 today and it delivers 3 gallons of oil and 2 bushels of corn tomorrow. If someone offers a new contract that delivers 12 gallons of oil and 11 bushels of corn tomorrow, what would be the price today of this new contract? A. $50 B. $52 C. $55 D. $57 E. $60

  2. Consider a put option written on the Apple Inc. stock that matures 1 month from now. Its strike price is $100 and the current price of Apple Inc. stock is $94. The option is traded at $5.5 today. What are the possible maximum and minimum net profit at maturity of a single short position in this put option, respectively?

    A. $5.5, −$94.5 B. $6, −$94
    C. $5.5, $0
    D. $94.5, −$5.5 E. $0, −$5.5

  3. A crude oil futures price expiring in 9 months from now is $150 while the spot price of crude oil is $145. Assume that the risk-free rate is 5% and storage yield of crude oil is 2%, both in continuously compounded annual terms. What is the implied convenience yield of crude oil in continuously compounded annual terms?

    A. 2.2% B. 2.5% C. 2.7% D. 2.8% E. 3%
Answer #1

(1) Contract A delivers 2 gallons of oil and 4 bushels of corn for $ 16 whereas Contract B delivers 3 gallons of oil and 2 bushels of corn for $ 12.

Let the price per gallon of oil be $ g and price per bushel of corn be $ c

Therefore, 2g+4c = 16 - (A) and 3g+2c = 12 - (B)

Multiplying equation (B) with 2 and subtracting the product from equation (A) we get:

6g + 4c - 2g - 4c = 24 - 16

4g = 8

g = $ 2 per gallon

Substituting g= $ 2 in equation (A) we get:

2 x 2 + 4c = 16

c = 12 / 4 = $ 3 per bushel

Therefore, price of contract offering 12 gallons of oil and 11 bushels of corn should be P = 12 x 2 + 11 x 3 = $ 57

Hence, the correct option is (D)

NOTE: Please raise separate queries for solutions to the remaining unrelated questions, as one query is restricted to the solution of only one complete question with a maximum of four sub-parts.